### v1

# A subcopula characterization of dependence for the Multivariate Bernoulli Distribution

Abstract
By applying Sklar’s theorem to the Multivariate Bernoulli Distribution (MBD), it is proposed a framework that decouples the marginal distributions from the dependence structure, providing a clearer understanding of how binary variables interact. Explicit formulas are derived under the MBD using subcopulas to introduce dependence measures for interactions of all orders, not just pairwise. A bayesian inference approach is also applied to estimate the parameters of the MBD, offering practical tools for parameter estimation and dependence analysis in real-world applications. The results obtained contribute to the application of subcopulas of multivariate binary data, with a real data example of comorbidities in COVID-19 patients.
Keywords: Sklar theorem, subcopula, dependence, Bernoulli, binary data.
1 Introduction
In statistical modeling, the analysis of multivariate binary data often requires flexible methods to capture the dependence structure between variables. The Multivariate Bernoulli Distribution (MBD) offers a natural framework for modeling such data, yet specifying and understanding the dependence between variables in this context can be challenging. This is especially true when working with multivariate data where traditional correlation measures fall short due to the binary nature of the variables and the dimensionality.
A representation of the MBD that provides an alternative to the traditional log-linear model for binary variables was proposed by [17] using the concept of Kronecker product from matrix calculus, in terms of $12^{n}-$ 1 parameters, where $n$ is the dimension of a random vector with MBD, but where the dependency vector of parameters are, in fact, non standardized central moments. In [11] several flexible methods for simulating random binary sequences with fixed marginal distributions and specified degrees of association between the variables are discussed, but only pairwise dependencies are considered. An algorithm is derived in [7] for generating systems of correlated binary data, but allowing for the specification of just pairwise correlations within each system. A simple method to characterize multivariate Bernoulli variables with given means is investigated in [5] but only pairwise are correlations are discussed.
Copulas, and more generally subcopulas, provide a straightforward approach to describing the dependence structure in multivariate settings by decoupling the marginal distributions from the joint distribution. Introduced by Sklar’s theorem [15], subcopulas extend copula theory to cases where the marginals are not necessarily continuous, making them particularly useful for multivariate discrete data. This paper focuses on leveraging subcopulas to characterize the dependence within the MBD, offering a framework to better understand and quantify dependence of all orders (pariwise, three-wise, four-wise, etc.)
Building upon previous work in copula theory and dependence measures, the present work introduces a formal subcopula-based characterization of the dependence structure within the MBD, and illustrates its usefulness to tackle the compatibility problem of building multivariate models in terms of lower dimension marginals.
This work provides a systematic approach for analyzing the joint dependence of multivariate binary data, with potential applications in fields ranging from genetics to econometrics where binary outcomes are prevalent. The results presented here not only contribute to the theoretical understanding of subcopulas but also offer practical tools for applied statistics and data analysis, and Julia programming [2] code is provided for their computational implementation and application.
2 Subcopulas and dependence
Sklar’s Theorem [15] proves that for any $n-$dimensional vector of random variables $(X_{1},\ldots,X_{n})$ there exists a functional link $S^{(n)}$ between its probability joint distribution $F_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\mathbb{P}(X_{1}\leq x_{1},\ldots,X_{n}\leq x_{n})$ and its corresponding univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i}),$ $i\in\{1,\ldots,n\}:$
$F_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})\,=\,S^{(n)}\left(F_{X_{1}}(x_{1}),\ldots,F_{X_{n}}(x_{n})\right)\,,$ (1)
for any $(x_{1},\ldots,x_{n})\in\overline{\mathbb{R}}^{n},$ and where the function $S^{(n)}:D_{1}\times\cdots\times D_{n}\rightarrow[0,1]$ is unique, with $D_{i}=\mbox{Ran}\,F_{X_{i}}$ for $i\in\{1,\ldots,n\},$ which implies $\{0,1\}\subseteq D_{i}\subseteq[0,1].$ Such link $S^{(n)}$ is called a subcopula function, and in the particular case where every $D_{i}=[0,1]$ then it is called a copula function, which would be the case when all the random variables are continuous. If at least one of the random variables is not continuous then the domain of the underlying subcopula $S^{(n)}$ is a proper subset of the unit hypercube $[0,1]^{n}.$ A formal definition for subcopula functions and their properties are discussed in detail in [13] and [12].
Recalling that the random variables $X_{1},\ldots,X_{n}$ are independent if and only if their probability joint distribution is equal to the product of its univariate marginals, then as an immediate consequence of (1) the unique subcopula function under independence is:
$\Pi^{(n)}(u_{1},\ldots,u_{n}):=u_{1}\cdots u_{n}\,,\qquad(u_{1},\ldots,u_{n})\in D_{1}\times\cdots\times D_{n}\,.$ (2)
Moreover, as a consequence of the Fréchet-Hoeffding bounds for joint distributions (see [6] and [8]) and applying (1) we may obtain bounds for any subcopula $S^{(n)}:$
$W^{(n)}(u_{1},\ldots,u_{n})\,\leq\,S^{(n)}(u_{1},\ldots,u_{n})\,\leq\,M^{(n)}(u_{1},\ldots,u_{n})\,,$ (3)
where $W^{(n)}(u_{1},\ldots,u_{n}):=\max\{u_{1}+\cdots+u_{n}-n+1,0\}$ and $M^{(n)}(u_{1},\ldots,u_{n}):=\min\{u_{1},\ldots,u_{n}\}.$ The upper bound $M^{(n)}$ is always a subcopula, but the lower bound $W^{(n)}$ is subcopula only for $n=2,$ though it is still the best possible lower bound for $n>2,$ see [12].
From any $n-$dimensional subcopula $S^{(n)}(u_{1},\ldots,u_{n})$ and $2\leq k<n$ it is possible to obtain $\binom{n}{k}$ $k-$dimensional marginal subcopulas $S^{(k)}(v_{1},\ldots,v_{k})$ defined in terms of $S^{(n)}(u_{1},\ldots,u_{n})$ where $k$ of its entries are equal to exactly one of each $\{v_{1},\ldots,v_{k}\}$ and the remaining $n-k$ are equal to $1$
Since the univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i})$ have no information about how each random variable $X_{i}$ interacts with others, then as a consequence of (1) all the information about the dependence among the random variables $X_{1},\ldots,X_{n}$ is contained in their unique underlying subcopula $S^{(n)},$ and therefore any attempt to measure degrees or intensity of dependence should extract information from $S^{(n)}.$ For the particular case of bivariate copulas (when both random variables are continuous) several copula-based measures have been studied, such as the concordance measures by Spearman [16] or Kendall [10], or dependence measures by Schweizer-Wolff [14] or Hoeffding [8]. For the case of proper subcopulas (that are not copulas) a subcopula dependence measure has been proposed by Erdely [3].
3 Bivariate Bernoulli
[…]
Accordingly to (1) the domain of the underlying subcopula in this case is the set $\mbox{Dom}\,S^{(2)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}.$ From the formal definition of bivariate subcopula (see [12], for example) it always satisfies $S^{(2)}(u,0)=0=S^{(2)}(0,v),$ $S^{(2)}(u,1)=u,$ and $S^{(2)}(1,v)=v,$ and therefore:
$S^{(2)}(u,v)=\begin{cases}\,1&\mbox{ if }(u,v)=(1,1),\\ \,\theta_{1}&\mbox{ if }(u,v)=(\theta_{1},1),\\ \,\theta_{2}&\mbox{ if }(u,v)=(1,\theta_{2}),\\ \,\theta_{12}&\mbox{ if }(u,v)=(\theta_{1},\theta_{2}),\\ \,0&\mbox{ elsewhere,}\end{cases}$ (4)
where $\theta_{12}$ is any value satisfying:
$\max\{\theta_{1}+\theta_{2}-1,0\}\,\leq\,\theta_{12}\,\leq\,\min\{\theta_{1},\theta_{2}\}$ (5)
as a consequence of (3), that in this case acts as a bivariate dependence parameter. Since $\mbox{Ran}\,(X,Y)=\{(0,0),(0,1),(1,0),(1,1)\}$ let $\{p_{00},p_{01},p_{10},p_{11}\}$ be the joint point probabilities:
$p_{ij}:=\mathbb{P}(X_{1}=i,X_{2}=j)\,,\qquad i,j\in\{0,1\},$ (6)
where necessarily $0\leq p_{ij}\leq 1$ and $p_{00}+p_{01}+p_{10}+p_{11}=1,$ and therefore only three of the $p_{ij}$ values need to be specified, since the fourth is just $1$ minus the sum of the other three. Applying (1) and (4) to calculate each $p_{ij}$ in terms of the joint distribution function $F_{12}(i,j)=\mathbb{P}(X_{1}\leq i,X_{2}\leq j):$
$p_{ij}$ $=$ $F_{12}(i,j)-F_{12}(i,j-1)-F_{12}(i-1,j)+F_{12}(i-1,j-1)\,,$ (7)
$=$ $S^{(2)}\left(F_{1}(i),F_{2}(j)\right)-S^{(2)}\left(F_{1}(i),F_{2}(j-1)\right)-S^{(2)}\left(F_{1}(i-1),F_{2}(j)\right)\ldots$
$+S^{(2)}\left(F_{1}(i-1),F_{2}(j-1)\right)\,,$
and therefore:
$p_{00}$ $=$ $F_{12}(0,0)=S^{(2)}(\theta_{1},\theta_{2})=\theta_{12}\,,$ (8)
$p_{01}$ $=$ $F_{12}(0,1)-F_{12}(0,0)=S^{(2)}(\theta_{1},1)-S^{(2)}(\theta_{1},\theta_{2})=\theta_{1}-\theta_{12}\,,$
$p_{10}$ $=$ $F_{12}(1,0)-F_{12}(0,0)=S^{(2)}(1,\theta_{2})-S^{(2)}(\theta_{1},\theta_{2})=\theta_{2}-\theta_{12}\,,$
$p_{11}$ $=$ $1-\theta_{1}-\theta_{2}+\theta_{12}\,.$
Formulas (8) allow to fully characterize the family $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ of all bivariate Bernoulli distributions in terms of the marginal parameters $10<\theta_{1}<$ 1 and $0<\theta_{2}<1,$ and a bivariate dependence parameter $\theta_{12}$ that must satisfy $\max\{\theta_{1}+\theta_{2}-1,0\}\leq\theta_{12}\leq\min\{\theta_{1},\theta_{2}\}.$ The subcopula based dependence measure proposed by [3] for $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ becomes:
$\mu(X_{1},X_{2})=\begin{cases}\frac{\theta_{12}-\theta_{1}\theta_{2}}{\min\{\theta_{1},\theta_{2}\}-\theta_{1}\theta_{2}}&\mbox{ if }\theta_{12}\geq\theta_{1}\theta_{2}\,,\\ {}&{}\textstyle\\ \frac{\theta_{12}-\theta_{1}\theta_{2}}{\theta_{1}\theta_{2}-\max\{\theta_{1}+\theta_{2}-1,0\}}&\mbox{ if }\theta_{12}<\theta_{1}\theta_{2}\,.\end{cases}$ (9)
Two Bernoulli random variables are independent if and only if $\theta_{12}=\theta_{1}\theta_{2}$ as an immediate consequence of (2), or equivalently if and only if $\mu(X_{1},X_{2})=0,$ and therefore:
$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1}\theta_{2},\theta_{1}(1-\theta_{2}),(1-\theta_{1})\theta_{2},(1-\theta_{1})(1-\theta_{2}))\,,$ (10)
as expected. For two dependent Bernoulli random variables, that is when $\theta_{12}\neq\theta_{1}\theta_{2},$ or equivalently when $\mu(X_{1},X_{2})\neq 0,$ without loss of generality assume that $\theta_{1}\leq\theta_{2}.$ The Fréchet-Hoeffding upper bound (3) is attained with dependence parameter $\theta_{12}=\min\{\theta_{1},\theta_{2}\}=\theta_{1}$ and therefore $\mu(X_{1},X_{2})=+1$ with:
$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1},0,\theta_{2}-\theta_{1},1-\theta_{2})\,.$ (11)
Only the particular case $\theta_{1}=\theta_{2}=\theta$ translates into:
$(p_{00},p_{01},p_{10},p_{11})=(\theta,0,0,1-\theta)\,,$ (12)
that is $\mathbb{P}(X_{1}=X_{2})=p_{00}+p_{11}=1.$ Equivalently, if $\theta_{1}\neq\theta_{2}$ then it is definitely not possible to get $\mathbb{P}(X_{1}=X_{2})=1.$
The Fréchet-Hoeffding lower bound (3) is reached with dependence parameter $\theta_{12}=\max\{\theta_{1}+\theta_{2}-1,0\}$ which applied to formulas (8) translates into:
$(p_{00},p_{01},p_{10},p_{11})=\begin{cases}\,(\theta_{1}+\theta_{2}-1,1-\theta_{2},1-\theta_{1},0)&\mbox{if }\theta_{1}+\theta_{2}>1\,,\\ \,(0,\theta_{1},\theta_{2},1-\theta_{1}-\theta_{2})&\mbox{if }\theta_{1}+\theta_{2}\leq 1\,,\end{cases}$ (13)
with $\mu(X_{1},X_{2})=-1,$ from where it is clear that $\mathbb{P}(X_{1}=1-X_{2})=1$ if and only if $\theta_{1}+\theta_{2}=1.$
In summary, formulas (8) allow to define the family $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ of all bivariate distributions with Bernoulli univariate marginals, which is in fact also known as a Fréchet class of the type $\mathcal{F}(F_{1},F_{2}),$ see for example [9]. The bivariate joint probability mass function $p_{ij}=\mathbb{P}(X_{1}=i,X_{2}=j)$ for $i$ and $j$ in $\{0,1\}$ can be specified in two ways:
  • a)
    choosing freely the two marginal parameters $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1 and then choosing $\theta_{12}$ such that (5) holds, or
  • b)
    choosing freely the dependence parameter $10<\theta_{12}<$ 1 and then choosing the marginal parameters $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1 such that (5) holds. Equivalently, instead of choosing the dependence parameter value, we may choose a given value of the dependence measure defined by [3] and then obtain the value for $\theta_{12}$ from (9).
Behind the above cases a) and b) is what is known as a compatibility problem, in this particular case for the Fréchet class $\mathcal{F}(F_{1},F_{2}):$ eventhough the three parameters are the probabilities $\theta_{1}=\mathbb{P}(X_{1}=0),$ $\theta_{2}=\mathbb{P}(X_{2}=0),$ and $\theta_{12}=\mathbb{P}(X_{1}=0,X_{2}=0),$ not all values in the $[0,1]$ interval are simultaneously admissible (or compatible) for such parameters, since (5) must hold, see Figure 1.
Figure 1: Bivariate Bernoulli. Left: Compatible values for the bivariate dependence parameter $\theta_{12}$ (red interval), given a specific pair of values $(\theta_{1}^{*},\theta_{2}^{*})$ for the univariate marginal parameters. Right: Compatible values for the marginal univariate parameters $(\theta_{1},\theta_{2})$ (red triangle region), given a specific value $\theta_{12}^{*}$ for the bivariate dependence parameter.
4 Trivariate Bernoulli
Let $(X_{1},X_{2},X_{3})$ be a vector of Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $10<\theta_{r}<$ 1 as explained in the previous section, but now $r\in\{1,2,3\}.$ Accordingly to (1) the domain of the underlying trivariate subcopula is the set $\mbox{Dom}\,S^{(3)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}\times\{0,\theta_{3},1\}$ and:
$S^{(3)}(u,v,w)=\begin{cases}\,1&\mbox{ if }(u,v,w)=(1,1,1),\\ \,\theta_{1}&\mbox{ if }(u,v,w)=(\theta_{1},1,1),\\ \,\theta_{2}&\mbox{ if }(u,v,w)=(1,\theta_{2},1),\\ \,\theta_{3}&\mbox{ if }(u,v,w)=(1,1,\theta_{3}),\\ \,\theta_{12}&\mbox{ if }(u,v,w)=(\theta_{1},\theta_{2},1),\\ \,\theta_{13}&\mbox{ if }(u,v,w)=(\theta_{1},1,\theta_{3}),\\ \,\theta_{23}&\mbox{ if }(u,v,w)=(1,\theta_{2},\theta_{3}),\\ \,\theta_{123}&\mbox{ if }(u,v,w)=(\theta_{1},\theta_{2},\theta_{3}),\\ \,0&\mbox{ elsewhere.}\end{cases}$ (14)
Recall from Section 2 that for a trivariate subcopula $S^{(3)}$ there are $\binom{3}{2}=3$ bivariate marginal subcopulas, which in this case have bivariate dependence parameters $\theta_{12},$ $\theta_{13},$ and $\theta_{23},$ respectively, plus a trivariate dependence parameter $\theta_{123}.$ The bivariate dependence parameters must satisfy (5), that is:
$\max\{\theta_{r}+\theta_{t}-1,0\}\,\leq\,\theta_{rt}\,\leq\,\min\{\theta_{r},\theta_{t}\}\,,\quad 1\leq r<t\leq 3\,.$ (15)
The trivariate dependence parameter $\theta_{123}$ must satisfy the Fréchet-Hoeffding bounds (3), which in this case translates into:
$\max\{\theta_{1}+\theta_{2}+\theta_{3}-2,0\}\,\leq\,\theta_{123}\,\leq\,\min\{\theta_{1},\theta_{2},\theta_{3}\}\,,$ (16)
but it will be discussed later that (16) is a necessary but could not be a sufficient condition for an admisible value for $\theta_{123},$ due to a compatibility problem, similarly as mentioned in the bivariate case. Since $\mbox{Ran}\,(X_{1},X_{2},X_{3})=\{(i,j,k):i,j,k\in\{0,1\}\}$ let:
$p_{ijk}:=\mathbb{P}(X_{1}=i,X_{2}=j,X_{3}=k)\,,\qquad i,j,k\in\{0,1\},$ (17)
where necessarily $0\leq p_{ijk}\leq 1$ and $\sum_{i,j,k}p_{i,j,k}=1,$ and therefore only seven of the $p_{ijk}$ need to be specified, since the eighth is just $1$ minus the sum of the other seven. Applying (1) and (14) to calculate each $p_{ijk}$ in terms of the joint distribution function $F_{123}(X_{1}\leq i,X_{2}\leq j,X_{3}\leq k):$
$p_{ijk}$ $=$ $F_{123}(i,j,k)-F_{123}(i,j,k-1)-F_{123}(i,j-1,k)\ldots$
$-F_{123}(i-1,j,k)+F_{123}(i,j-1,k-1)+F_{123}(i-1,j,k-1)\ldots$
$+F_{123}(i-1,j-1,k)-F_{123}(i-1,j-1,k-1)\,,$
$=$ $S^{(3)}\left(F_{1}(i),F_{2}(j),F_{3}(k)\right)-S^{(3)}\left(F_{1}(i),F_{2}(j),F_{3}(k-1)\right)\ldots$
$-S^{(3)}\left(F_{1}(i),F_{2}(j-1),F_{3}(k)\right)-S^{(3)}\left(F_{1}(i-1),F_{2}(j),F_{3}(k)\right)\ldots$
$+S^{(3)}\left(F_{1}(i),F_{2}(j-1),F_{3}(k-1)\right)+S^{(3)}\left(F_{1}(i-1),F_{2}(j),F_{3}(k-1)\right)\ldots$
$+S^{(3)}\left(F_{1}(i-1),F_{2}(j-1),F_{3}(k)\right)-S^{(3)}\left(F_{1}(i-1),F_{2}(j-1),F_{3}(k-1)\right)\,,$
and therefore:
$p_{000}$ $=$ $\theta_{123}\,,$ (19)
$p_{001}$ $=$ $\theta_{12}-\theta_{123}\,,$ (20)
$p_{010}$ $=$ $\theta_{13}-\theta_{123}\,,$ (21)
$p_{100}$ $=$ $\theta_{23}-\theta_{123}\,,$ (22)
$p_{011}$ $=$ $\theta_{1}-\theta_{12}-\theta_{13}+\theta_{123}\,,$ (23)
$p_{101}$ $=$ $\theta_{2}-\theta_{12}-\theta_{23}+\theta_{123}\,,$ (24)
$p_{110}$ $=$ $\theta_{3}-\theta_{13}-\theta_{23}+\theta_{123}\,,$ (25)
$p_{111}$ $=$ $1-\theta_{1}-\theta_{2}-\theta_{3}+\theta_{12}+\theta_{13}+\theta_{23}-\theta_{123}\,,$ (26)
where (26) is equal to $1$ minus the sum of (19) through (25). From (20), (21), (22), and (16), it is clear that:
$\theta_{123}\,\leq\,\min\{\theta_{12},\theta_{13},\theta_{23}\}\,\leq\,\min\{\theta_{1},\theta_{2},\theta_{3}\}$ (27)
a sharper upper bound for $\theta_{123}$ than the Fréchet-Hoeffding bounds (16), as a consequence of the compatibility problem for the Fréchet class $\mathcal{F}\{F_{12},F_{13},F_{23}\},$ but all the inequalities that arise from (19) to (26) have to be analyzed under the restrictions $0\leq p_{ijk}\leq 1$ to solve the full compatibility problem of all the parameters. It is straightforward to obtain expressions for the $\theta$ parameters in terms of the $p_{ijk}$ probabilities:
$\theta_{123}$ $=$ $p_{000}=\mathbb{P}(X_{1}=0,X_{2}=0,X_{3}=0)\,,$ (28)
$\theta_{12}$ $=$ $p_{000}+p_{001}=\mathbb{P}(X_{1}=0,X_{2}=0)\,,$ (29)
$\theta_{13}$ $=$ $p_{000}+p_{010}=\mathbb{P}(X_{1}=0,X_{3}=0)\,,$ (30)
$\theta_{23}$ $=$ $p_{000}+p_{100}=\mathbb{P}(X_{2}=0,X_{3}=0)\,,$ (31)
$\theta_{1}$ $=$ $p_{000}+p_{001}+p_{010}+p_{011}=\mathbb{P}(X_{1}=0)\,,$ (32)
$\theta_{2}$ $=$ $p_{000}+p_{001}+p_{100}+p_{101}=\mathbb{P}(X_{2}=0)\,,$ (33)
$\theta_{3}$ $=$ $p_{000}+p_{010}+p_{100}+p_{110}=\mathbb{P}(X_{3}=0)\,.$ (34)
We may follow the order in (28) through (34) to specify all the parameters such that the full compatibility of the Fréchet class $\mathcal{F}\{F_{12},F_{13},F_{23},F_{1},F_{2},F_{3}\}$ is guaranteed:
Algorithm 1
  • Step 0:
    Choose $0\leq p_{000}\leq 1$ and set the trivariate dependence parameter $\theta_{123}=p_{000}\,.$
  • Step 1:
    Choose $0\leq p_{001}\leq 1,$ $0\leq p_{010}\leq 1,$ and $0\leq p_{100}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}\leq 1,$ and calculate the bivariate dependence parameters $\theta_{rt}$ accordingly to (29), (30), and (31).
  • Step 2:
    Choose $0\leq p_{011}\leq 1,$ $0\leq p_{101}\leq 1,$ and $0\leq p_{110}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}+p_{011}+p_{101}+p_{110}\leq 1,$ and calculate the univariate (marginal) parameters $\theta_{r}$ accordingly to (32), (33), and (34).
  • Step 3:
    Calculate $p_{111}=1-p_{000}-p_{001}-p_{010}-p_{011}-p_{100}-p_{101}-p_{110}\,.$
Finally, we may apply (9) to calculate the bivariate dependence measures $\mu_{rt}=\mu(X_{r},X_{t})$ for $1\leq r<t\leq 3,$ and following the same idea:
$\mu_{123}=\begin{cases}\frac{\theta_{123}-\theta_{1}\theta_{2}\theta_{3}}{\min\{\theta_{1},\theta_{2},\theta_{3}\}-\theta_{1}\theta_{2}\theta_{3}}&\mbox{ if}\theta_{123}\geq\theta_{1}\theta_{2}\theta_{3}\,,\\ {}&{}\textstyle\\ \frac{\theta_{123}-\theta_{1}\theta_{2}\theta_{3}}{\theta_{1}\theta_{2}\theta_{3}-\max\{\theta_{1}+\theta_{2}+\theta_{3}-2,0\}}&\mbox{ if }\theta_{123}<\theta_{1}\theta_{2}\theta_{3}\,.\end{cases}$ (35)
where from (16) $-1\leq\mu_{123}\leq+1,$ and clearly $\mu_{123}=0$ if and only if the three random variables are independent.
Example 1. Consider $X_{1},X_{2},X_{3}$ identically distributed Bernoulli random variables with parameter $1-\theta,$ where $0<\theta<1,$ that are pairwise independent but not necessarily jointly independent. We will obtain all the compatible values for the trivariate dependence parameter $\theta_{123},$ and the resulting trivariate probability mass function (pmf). Combining the inequalities $0\leq p_{ijk}\leq 1$ (19) through (26) with $\theta_{r}=\theta$ and $\theta_{rt}=\theta^{2}$ (pairwise independence) we obtain:
$\max\{0,2\theta^{2}-\theta\}\,\leq\,\theta_{123}\,\leq\,\min\{\theta^{2},1+\theta-2\theta^{2},1-3\theta+3\theta^{2}\}$ (36)
which is equivalent to:
$\theta_{123}\in\begin{cases}\,\left[0\,,\,\theta^{2}\right]&\text{ if }\theta\leq\frac{1}{2}\\ \,\left[2\theta^{2}-\theta\,,\,1-3\theta+3\theta^{2}\right]&\text{ if }\frac{1}{2}<\theta\leq\frac{4}{5}\\ \,\left[2\theta^{2}-\theta\,,\,1+\theta-2\theta^{2}\right]&\text{ if }\frac{4}{5}<\theta\leq\frac{1+\sqrt{5}}{4}\end{cases}$ (37)
In Figure 2 it is depicted the set of all compatibility values $(\theta,\theta_{123}),$ and in Table 1 the resulting pmf in the particular case when $\theta=1/2$ and the special cases of extreme dependence ($\theta_{123}\in\{0,1/4\}$) and independence ($\theta_{123}=1/8$).
Figure 2: Compatible values $(\theta,\theta_{123})$ (red region)
in Example 1 for a Trivariate Bernoulli, with identical univariate marginals with parameter $1-\theta,$ pairwise independent, with trivariate dependence parameter $\theta_{123}.$
$\mu_{123}=-1$ $\mu_{123}=0$ $\mu_{123}=+1/3$
$(i,j,k)$ $p_{ijk}$ $\theta_{123}=0$ $\theta_{123}=1/8$ $\theta_{123}=1/4$
$(0,0,0)$ $\theta_{123}$ 00 $1/8$ $1/4$
$(0,0,1)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
$(0,1,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
$(1,0,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
$(0,1,1)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
$(1,0,1)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
$(1,1,0)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
$(1,1,1)$ $1-3\theta+3\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
Table 1: Probability mass function of a trivariate Bernoulli in Example 1, with identical univariate marginals with parameter $1-\theta=1/2,$ pairwise independent, and with trivariate parameter $\theta_{123}\in\{0,1/8,1/4\}$ corresponding to the lowest compatible value, independence, and the highest compatible value, respectively.
Example 2. Algorithm 1 will be applied to obtain non independent $X_{1},X_{2},X_{3}$ but identically distributed Bernoulli random variables with parameter $1-\theta,$ where $0<\theta<1,$ with pairwise conditional independence, but not pairwise independent. For the sake of simplicity we will assume that the bivariate dependence parameter is the same for all pairs, say $\beta:=\theta_{12}=\theta_{13}=\theta_{23}$ and let $\tau:=\theta_{123}$ be the trivariate dependence parameter.
  • Step 1)
    Choose $p_{000}\in[0,1]$ and set the trivariate dependence parameter $\tau=p_{000}.$
  • Step 2)
    Choose $p_{001},p_{010},p_{100}\in[0,1]$ such that:
    $p_{000}+p_{001}+p_{010}+p_{100}\leq 1$ (38)
    but from (20), (21), and (22), we get $p_{001}=\beta-\tau=p_{010}=p_{100}$ and therefore $\beta\geq\tau.$ From (38) $p_{001}+p_{010}+p_{100}\leq 1-\tau$ implies $\beta\leq(1+2\tau)/3,$ so combining altogether:
    $\tau\,\leq\,\beta\,\leq\,\frac{1+2\tau}{3}$ (39)
  • Step 3)
    Choose $p_{011},p_{101},p_{110}\in[0,1]$ such that:
    $p_{000}+p_{001}+p_{010}+p_{100}+p_{011}+p_{101}+p_{110}\leq 1$ (40)
    but from (23), (24), and (25), we get $p_{011}=\theta-2\beta+\tau=p_{101}=p_{110}$ and therefore from (40) we obtain:
    $\theta\,\leq\,\frac{1-\tau}{3}+\beta$ (41)
  • Step 4)
    Calculate $p_{111}=1-\tau-3(\beta-\tau)-3(\theta-2\beta+\tau)=1-\tau+3(\beta-\theta).$
In summary, once specified the trivariate dependence parameter $0\leq\tau\leq 1$ then we have to choose the bivariate dependence parameter $\beta$ according to (39) and the univariate parameter $\theta$ according to (41), for compatibility. If in addition we want pairwise conditional independence, say for example $X_{1}$ and $X_{2}$ conditionally independent given the value of $X_{3}:$
$X_{1}\perp X_{2}\,|\,X_{3}=k\,\,\text{ where }k\in\{0,1\}\quad\Leftrightarrow\\ \begin{cases}\mathbb{P}(X_{1}=0,X_{2}=0\,|\,X_{3}=0)=\mathbb{P}(X_{1}=0\,|\,X_{3}=0)\mathbb{P}(X_{2}=0\,|\,X_{3}=0)\\ \mathbb{P}(X_{1}=0,X_{2}=0\,|\,X_{3}=1)=\mathbb{P}(X_{1}=0\,|\,X_{3}=1)\mathbb{P}(X_{2}=0\,|\,X_{3}=1)\end{cases}$
$X_{1}\perp X_{2}\,|\,X_{3}=k$$\Leftrightarrow$ $\begin{cases}\frac{p_{000}}{\theta}=\frac{p_{000}+p_{010}}{\theta}\cdot\frac{p_{000}+p_{100}}{\theta}\\ \frac{p_{001}}{1-\theta}=\frac{p_{001}+p_{011}}{1-\theta}\cdot\frac{p_{001}+p_{101}}{1-\theta}\end{cases}$ (42)
$\Leftrightarrow$ $\begin{cases}\theta\tau=\beta^{2}\\ (1-\theta)(\beta-\theta)=(\theta-\beta)^{2}\end{cases}$
$\Leftrightarrow$ $\tau^{2}(\beta-\tau)=\beta^{3}(\beta-\tau)$
where (42) is always satisfied whenever $\beta=\tau$ which in turn would imply that $\theta=\beta$ and since $\beta\neq\theta^{2}$ nor $\tau=\theta^{3}$ for $10<\theta<$ 1 then the condition $\theta=\beta=\tau$ implies pairwise dependence and trivariate dependence but with pairwise conditional independence. In fact, in terms of the probabilities $p_{ijk}$ this implies $p_{000}=\theta,$ $p_{111}=1-\theta$ and zero probability for the remaining six cases.
5 Multivariate Bernoulli
Let $(X_{1},\ldots,X_{n})$ be a vector of $n\geq 2$ Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1,$ $r\in\{1,\ldots,n\}.$ Accordingly to (1) the domain of the underlying multivariate subcopula is the set $\operatorname{Dom}\,S^{(n)}=\{0,\theta_{1},1\}\,\times\,\cdots\,\times\,\{0,\theta_{n},1\}$ and:
$S^{(n)}(u_{1},\ldots,u_{n})=\begin{cases}\,1&\mbox{ if }u_{r}=1\mbox{ for all }r\in\{1,\ldots,n\},\\ \,\theta_{s}&\mbox{ if }u_{s}=\theta_{s}\mbox{ and }u_{r}=1\mbox{ for all }r\neq s,\\ \,\theta_{rs}&\mbox{ if }r<s,u_{r}=\theta_{r},u_{s}=\theta_{s},\mbox{ and }u_{t}=1\mbox{ for all }t\neq r,s,\\ \,\theta_{rst}&\mbox{ if }r<s<t,u_{r}=\theta_{r},u_{s}=\theta_{s},u_{t}=\theta_{t},u_{\ell}=1,\ell\neq r,s,t,\\ \quad\vdots&\qquad\vdots\\ \,\theta_{12\cdots n}&\mbox{ if }u_{1}=\theta_{1},\ldots,u_{n}=\theta_{n},\\ \,0&\mbox{ elsewhere.}\end{cases}$ (43)
with a total of $12^{n}-$ 1 parameters. For any ordered subset of $\{1,\ldots,n\},$ say $A=\{a_{1},\ldots,a_{m}\}$ where $m\leq n$ and if $r<s$ then $a_{r}<a_{s},$ the $m$-variate dependence parameter $\theta_{\!A}$ must satisfy the Fréchet-Hoeffding bounds (3), which in this case translates into:
[…]
Since $\mbox{Ran}\,(X_{1},\ldots,X_{n})=\{0,1\}^{n}$ let:
$p_{i_{1}\ldots i_{n}}:=\mathbb{P}(X_{1}=i_{1},\ldots,X_{n}=i_{n})\,,\qquad i_{r}\in\{0,1\},$ (45)
where necessarily $0\leq p_{i_{1}\ldots i_{n}}\leq 1$ and $\sum\cdots\sum p_{i_{1}\ldots i_{n}}=1,$ and therefore just $12^{n}-$ 1 of the $p_{i_{1}\ldots i_{n}}$ need to be specified, since the $2^{n}$-th is just $1$ minus the sum of the other $12^{n}-$ 1 . Then expressions for the $\theta_{\!A}$ parameters in terms of (45) probabilities are obtained by:
$\theta_{\!A}\,=\,\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\qquad\qquad i_{r}\,\notin\,A}\!\!\!\!\!\!\!\!\!\!\!\!\cdots\sum p_{i_{1}\ldots i_{n}}$ (46)
which can be specified according to the following algorithm to guarantee full compatibility of the dependence parameters, with the conventional notation that if $D$ is a set of numbers then $\sum D$ is equivalent to $\sum_{d\in D}d\,:$
Algorithm 2
  • Step 0:
    Choose $0\leq p_{0\cdots 0}\leq 1,$ and define the singleton set $B_{0}:=\{p_{0\cdots 0}\}.$ Set the $n$-variate dependence parameter $\theta_{1\cdots n}=p_{0\cdots 0}.$
  • Step 1:
    Choose $B_{1}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=1\}$ such that $\sum B_{0}\cup B_{1}\leq 1,$ and set the $(n-1)$-variate parameters $\theta_{\!A}$ accordingly to (46).
  • Step 2:
    Choose $B_{2}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=2\}$ such that $\sum B_{0}\cup B_{1}\cup B_{2}\leq 1,$ and set the $(n-2)$-variate parameters $\theta_{\!A}$ accordingly to (46).
  • Step $k$:
[…]
  • Step $n$:
[…]
Finally, we may calculate $m$-variate dependence measures for $2\leq m\leq n$ as follows:
$\mu_{A}=\begin{cases}\frac{\theta_{\!A}-\prod_{r\in A}\theta_{r}}{\min\{\theta_{a}:a\in A\}-\prod_{r\in A}\theta_{r}}&\mbox{ if }\theta_{\!A}\geq\prod_{r\in A}\theta_{r}\,,\\ {}&{}\textstyle\\ \frac{\theta_{\!A}-\prod_{r\in A}\theta_{r}}{\prod_{r\in A}\theta_{r}-\max\{\sum_{a\in A}\theta_{a}-m+1,0\}}&\mbox{ if }\theta_{\!A}<\prod_{r\in A}\theta_{r}\,.\end{cases}$ (47)
where from (44) $-1\leq\mu_{A}\leq+1,$ and clearly $\mu_{A}=0$ if and only if the random variables indexed by $A$ are independent.
5.1 Bayesian inference
If $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\}$ is a size $m$ observed random sample from a Multivariate Bernoulli random vector $\mathbf{X}=(X_{1},\ldots,X_{n})$ where $\mathbf{x}_{j}=(x_{j,1},\ldots,x_{j,n})$ for $j\in\{1,\ldots,m\},$ we can make statistical inferences about the $2^{n}$ parameters $p_{i_{1}\ldots i_{n}}.$ For a more simplified notation, following [17], since $(i_{1},\ldots,i_{n})\in\{0,1\}^{n}$ we may consider it as a $n$-positions binary number and apply the one-to-one mapping conversion to a positive integer. For $r\in\{1,\ldots,2^{n}-1\}$ let $B(r-1)=[i_{1}(r),\ldots,i_{n}(r)]$ be the $n$-positions binary representation of $r-1$, so that:
$p_{r}=\mathbb{P}[X_{1}=i_{1}(r),\ldots,X_{n}=i_{n}(r)]$ (48)
and:
[…]
Then the likelihood function associated to the unknown parameters $p_{1},\ldots,p_{2^{n}-1}$ is given by:
[…]
where:
[…]
Clearly (50) has as conjugate prior distribution a $12^{n}-$ 1 dimensional Dirichlet (Di) distribution with $2^{n}$ hyperparameters $\alpha_{r}>0,$ see for example [1], and therefore the posterior distribution is given by:
$\pi(\mathbf{p}\,|\,\mathbf{x})\sim\text{Di}_{\,2^{n}-1}(\alpha_{1}+n_{1},\ldots,\alpha_{2^{n}-1}+n_{2^{n}-1},\alpha_{2^{n}}+m-\sum_{r\,=\,1}^{2^{n}-1}n_{r})$ (52)
where $\mathbf{p}=(p_{1},\ldots,p_{2^{n}-1})$ and $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{m}).$ With (52) we can also make posterior inferences for the $\theta_{\!A}$ parameters through (46), and for the $\mu_{A}$ parameters through (47).
5.2 Computational implementation
The Julia programming language [2] has been used for the implementation of all the previous results, and the code is available in a public repository at [4] for full reproducibility. The main functions implemented:
  • MBerDep Calculates dependence parameters $\theta_{A}$ and measures $\mu_{A}$ given the parameters $p_{1},\ldots,p_{2^{n}}$ of a $n$-dimensional multivariate Bernoulli distribution.
  • MBerMargin Calculates dependence parameters and measures for the $m$-dimensional marginal distribution of a $n$-dimensional multivariate Bernoulli distribution, where $1\leq m\leq n.$
  • MBerCond Calculates conditional probabilities, dependence parameters and measures from a $n$-dimensional multivariate Bernoulli distribution, given the values for a subset of random variables.
  • MBerSim Simulates a random sample from a $n$-dimensional multivariate Bernoulli distribution.
  • MBerBayes Posterior Dirichlet model for the parameters of a $n$-dimensional multivariate Bernoulli distribution given an observed random sample.
  • MBerInf Posterior point and interval estimation for the probability parameters, and the dependence parameters and measures, given an observed sample from a multivariate Bernoulli distribution, and a prior value. For dimension $10$ or higher calculations may take more than an hour. In such case you may consider using MBerEst just for point estimations.
Example 1 revisited. If we choose $\theta=0.6$ then according to (37) we must choose $\theta_{123}\in[0.12,0.28]$ where $\theta^{3}=0.216$ belongs to such interval and would represent the case of joint independence. Let’s use $\theta_{123}=0.15$ and applying formulas (19) through (26) we may obtain the $p_{ijk}$ joint probabilities $p_{000}=0.15,$ $p_{001}=0.21,$ $p_{010}=0.21,$ $p_{011}=0.03,$ $p_{100}=0.21,$ $p_{101}=0.03,$ $p_{110}=0.03,$ and $p_{111}=0.13.$ With such values me define in Julia the vector:
julia> pp = [0.15,0.21,0.21,0.03,0.21,0.03,0.03,0.13];
julia> X = MBerDep(pp);
And then verify the we get the expected dependence parameters and measures:
julia> [X.dparam.idx X.dparam.value]
72 Matrix{Any}:
 [1, 2, 3]  0.15
 [1, 2]     0.36
 [1, 3]     0.36
 [2, 3]     0.36
 [1]        0.6
 [2]        0.6
 [3]        0.6
from where we get that the bivariate dependence parameters $\theta_{12}=0.36=\theta_{13}=\theta_{23}=\theta^{2}$ as expected (pairwise independence) but certainly not joint independence since $\theta_{123}=0.15\neq 0.216=\theta^{3}.$ For the dependence measures we get:
julia> [X.dmeas.idx X.dmeas.value]
42 Matrix{Any}:
 [1, 2, 3]  -0.305556
 [1, 2]      0.0
 [1, 3]      0.0
 [2, 3]      0.0
with pairwise dependencies $\mu_{ij}$ all equal to zero (as expected) but with a 3-variate negative dependence $\mu_{123}=-0.306$ as expected, since we chose a value $\theta_{123}=0.15$ below the value that would represent independence ($\theta^{3}=0.216$).
Example 3. A trivariate Bernoulli distribution will be defined applying Algorithm 1 from Section 4 as follows:
julia> begin
           p000 = 0.1
           p001, p010, p100 = 0.2, 0.1, 0.05
           p011, p101, p110 = 0.2, 0.15, 0.1
           p111 = 1 - sum([p000, p001, p010, p011, p100, p101, p110])
           pp = [p000, p001, p010, p011, p100, p101, p110, p111]
       end
8-element Vector{Float64}:
 0.1
 0.2
 0.1
 0.2
 0.05
 0.15
 0.1
 0.09999999999999987
julia> X = MBerDep(pp); X.nrvs
3
where X.nrvs is the corresponding dimension of the random vector. Then we simulate a random sample of size $3,000$ and make point and interval estimations about the probability parameters, and the dependence parameters and measures, and a comparison to the theoretical values. In the estimation matrices the third column is always a median posterior point estimation, and the second and fourth columns are the extreme points of a $99\%$ probability posterior interval estimation:
julia> begin
           Random.seed!(1234) # for reproducibility
           simX = MBerSim(pp, 3_000);
           infer = MBerInf(simX, prior = 1/2, nsim = 10_000, probint = 0.99);
           infer.probs # estimated probabilities
       end
84 Matrix{Any}:
 [0, 0, 0]  0.0861693  0.0996407  0.1143
 [0, 0, 1]  0.176657   0.19481    0.21376
 [0, 1, 0]  0.087351   0.100971   0.115828
 [0, 1, 1]  0.175131   0.193508   0.212438
 [1, 0, 0]  0.0332592  0.0420294  0.0520198
 [1, 0, 1]  0.139136   0.155535   0.173265
 [1, 1, 0]  0.090522   0.104274   0.119274
 [1, 1, 1]  0.0945739  0.108592   0.1238
julia> [X.binprob.idx X.binprob.value] # theoretical probabilities
82 Matrix{Any}:
 [0, 0, 0]  0.1
 [0, 0, 1]  0.2
 [0, 1, 0]  0.1
 [0, 1, 1]  0.2
 [1, 0, 0]  0.05
 [1, 0, 1]  0.15
 [1, 1, 0]  0.1
 [1, 1, 1]  0.1
julia> infer.dparam # estimated dependence parameters
74 Matrix{Any}:
 [1, 2, 3]  0.0861693  0.0996407  0.1143
 [1, 2]     0.273408   0.294522   0.316158
 [1, 3]     0.182337   0.200692   0.219919
 [2, 3]     0.125958   0.141746   0.158641
 [1]        0.566172   0.589215   0.612288
 [2]        0.468764   0.492305   0.515974
 [3]        0.325026   0.347202   0.369748
julia> [X.dparam.idx X.dparam.value] # theoretical dependence parameters
72 Matrix{Any}:
 [1, 2, 3]  0.1
 [1, 2]     0.3
 [1, 3]     0.2
 [2, 3]     0.15
 [1]        0.6
 [2]        0.5
 [3]        0.35
julia> infer.dmeas # estimated dependence measures
44 Matrix{Any}:
 [1, 2, 3]  -0.113144   -0.0104273   0.039112
 [1, 2]     -0.0338129   0.0222768   0.0793568
 [1, 3]     -0.072454   -0.0188205   0.0504544
 [2, 3]     -0.23491    -0.170452   -0.105142
julia> [X.dmeas.idx X.dmeas.value] # theoretical dependence measures
42 Matrix{Any}:
 [1, 2, 3]  -0.047619
 [1, 2]      0.0
 [1, 3]     -0.047619
 [2, 3]     -0.142857
Example 4. Data from official statistics about COVID-19 in Mexico during the first year of the pandemic, available for download from [4] as a covid2020.csv file, with 2.15 million patient registries and 15 binary variables measured (among many others available).
begin
    df = CSV.read("covid2020.csv", DataFrame)
    show(describe(df), allrows = true)
    data = zeros(Int, size(df))
    for c in 1:ncol(df)
        data[:, c] = df[:, c]
    end
    println()
    data
end
15x7 DataFrame
 Row | variable        mean        min    median   max    nmissing  eltype
     | Symbol          Float64     Int64  Float64  Int64  Int64     DataType
   1 | SEXO            0.501803        0      1.0      1         0  Int64
   2 | TIPO_PACIENTE   0.196705        0      0.0      1         0  Int64
   3 | DIABETES        0.1348          0      0.0      1         0  Int64
   4 | EPOC            0.0113235       0      0.0      1         0  Int64
   5 | INMUSUPR        0.00843405      0      0.0      1         0  Int64
   6 | HIPERTENSION    0.174666        0      0.0      1         0  Int64
   7 | CARDIOVASCULAR  0.0158313       0      0.0      1         0  Int64
   8 | OBESIDAD        0.144864        0      0.0      1         0  Int64
   9 | RENAL_CRONICA   0.0152094       0      0.0      1         0  Int64
  10 | TABAQUISMO      0.0735159       0      0.0      1         0  Int64
  11 | e00             0.0502618       0      0.0      1         0  Int64
  12 | e20             0.38682         0      0.0      1         0  Int64
  13 | e40             0.434831        0      0.0      1         0  Int64
  14 | e65             0.128087        0      0.0      1         0  Int64
  15 | MUERTE          0.101051        0      0.0      1         0  Int64
2154719x15 Matrix{Int64}:
 0  0  1  0  0  0  0  0  0  0  0  0  1  0  0
 1  0  0  0  0  0  0  0  0  0  0  1  0  0  0
 0  0  1  0  0  0  0  0  0  0  0  0  1  0  0
 1  1  1  1  0  0  1  1  0  0  0  0  0  1  1
 0  1  1  0  0  1  0  1  0  0  0  0  1  0  1
 1  1  0  0  0  0  0  1  0  0  0  0  0  1  1
With the following description of the variables:
  1. 1.
    SEXO = $1$ for male, 00 for female.
  2. 2.
    TIPO_PACIENTE = $1$ for hospitalized, 00 non-hospitalized.
  3. 3.
    DIABETES = $1$ for diabetic, 00 if not.
  4. 4.
    EPOC = $1$ for chronic obstructive pulmonary disease (COPD), 00 if not.
  5. 5.
    INMUSUPR = $1$ for immunosuppression, 00 if not.
  6. 6.
    HIPERTENSION = $1$ for hypertension, 00 if not.
  7. 7.
    CARDIOVASCULAR = $1$ for cardiovascular disease, 00 if not.
  8. 8.
    OBESIDAD = $1$ for obesity, 00 if not.
  9. 9.
    RENAL_CRONICA = $1$ for chronic kidney disease (CKD), 00 if not.
  10. 10.
    TABAQUISMO = $1$ for smoker, 00 for non-smoker.
  11. 11.
    e00 = $1$ for ages between 00 and $19$, 00 if not.
  12. 12.
    e20 = $1$ for ages between $20$ and $39$, 00 if not.
  13. 13.
    e40 = $1$ for ages between $40$ and $64$, 00 if not.
  14. 14.
    e65 = $1$ for ages between $65$ and older, 00 if not.
  15. 15.
    MUERTE = $1$ for death, 00 if survived.
In this case $n=15$ and therefore there are $7672^{15}-1=32,$ 767 probabilities to estimate, $12^{15}-$ 1 parameters to estimate, and $162^{15}-$ 16 multivariate dependencies to estimate (pairwise, three-wise, up to 15-wise).
@time estim = MBerEst(data); # 5 minutes approx
If we are interested in the 20 largest values for dependence parameters that include variable 15 (death) then:
begin
    iord = sortperm(estim.dmeas.value, rev = true)
    midx = estim.dmeas.idx[iord]
    m = estim.dmeas.value[iord]
    iMue = findall(x -> 15 in x, midx)
    [midx[iMue] m[iMue]][1:20, :]
end
202 Matrix{Any}:
 [2, 15]              0.902144
 [2, 4, 15]           0.859008
 [2, 9, 15]           0.858144
 [2, 5, 15]           0.855051
 [2, 4, 9, 15]        0.831122
 [2, 7, 15]           0.831098
 [2, 5, 9, 15]        0.828303
 [2, 4, 5, 15]        0.827159
 [2, 7, 9, 15]        0.809365
 [2, 4, 5, 9, 15]     0.808968
 [2, 4, 7, 15]        0.808805
 [2, 5, 7, 15]        0.803182
 [2, 4, 7, 9, 15]     0.793962
 [2, 5, 7, 9, 15]     0.789702
 [2, 4, 5, 7, 15]     0.787861
 [2, 4, 5, 7, 9, 15]  0.778416
 [2, 4, 9, 14, 15]    0.659362
 [2, 4, 14, 15]       0.658259
 [2, 14, 15]          0.656512
 [2, 9, 14, 15]       0.65601
Therefore from the 20 highest values for multivariate dependence that always include death we are able to identify in all cases variable 2 (hospitalization), and in several cases variables 4, 9, 5, 7 and 14 (COPD, CKD, immunosuppression, cardiovascular disease, and age 65 or older). In this data the mortality rate was $10.1\%$ but conditioning on the mentioned variables it increases up to $66.3\%:$
begin
    pd = mean(data[:, 15])
    estim2 = MBerEst(data[:, [2,4,5,7,9,14,15]])
    pdcond = MBerCond(estim2.binprob.value, [7], [1,2,3,4,5,6],
                      [1,1,1,1,1,1]).binprob.dic[[1]]
    println("P(death) = ", pd)
    println("P(death|hospital, COPD, immunosup, cardio, CKD,
             age 65+) = ", pdcond)
end
P(death) = 0.10105122756145929
P(death|hosp,COPD,immunosup,cardio,CKD,65+) = 0.6632653061224489
6 Conclusions
In this paper, a novel subcopula-based characterization of dependence for the Multivariate Bernoulli Distribution (MBD) has been introduced. By leveraging Sklar’s theorem, it has been shown that subcopulas can effectively capture the dependence structure between Bernoulli random variables, extending the utility of subcopula theory to binary data. The proposed approach provides a flexible framework for studying multivariate binary data by separating the marginal distributions from the joint dependence structure.
Explicit formulas and bounds for both bivariate and multivariate Bernoulli distributions in terms of subcopulas were derived, allowing for a more granular understanding of the dependence parameters. In particular, dependence measures were obtained for bivariate and trivariate Bernoulli distributions, and proposed a general method to extend these results to higher dimensions. This work’s findings confirm that subcopulas are a powerful tool for modeling and quantifying dependence in the MBD of all orders, not just pairwise.
The proposed methodology not only advances the theoretical understanding of subcopulas but also opens up new possibilities for practical applications in fields dealing with multivariate binary data. Potential applications include genetics, epidemiology, and social sciences, where understanding the dependence structure between binary variables is essential. Moreover, the bayesian inference framework presented offers a robust tool for parameter estimation and uncertainty quantification, enhancing the practical implementation of the proposed approach.
Computational implementations of the MBD are quite scarce, so the set of functions provided in subsection 5.2 and downloadable from [4] is a contribution that may help in analyzing and simulating multivariate binary data. Future research could explore further generalizations of the subcopula approach to other types of discrete data and investigate its performance in applied contexts. Additionally, incorporating covariates or other forms of heterogeneity into the subcopula framework may provide more nuanced models, particularly for data sets with more complex dependence structures.